OpenStudy (anonymous):

What is the exact area of a circle with a diameter of 15 inches?

OpenStudy (campbell_st):

do you know the formula for the area of a circle...?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Formula is pi times radius squared, it would become 176.625

OpenStudy (campbell_st):

ok... here is a good site that will help you http://www.mathopenref.com/ use this to help gain an understanding of some maths concepts

OpenStudy (anonymous):

Do I multiply 176.625 into 15?

OpenStudy (campbell_st):

lol.... wow... go and please read about some of the ideas you are asking about...

OpenStudy (anonymous):

Can you go over with me step by step?

OpenStudy (campbell_st):

nope.... best method is to go and read then ask about what you don't understand.. good luck

OpenStudy (anonymous):

Circumference is 2(pi)(r). r=(1/2)d=7.5 Now plug it in. C=(2)(7.5)pi=15pi=47.123889